The measure of an angle is 3 more than twice the measure of its supplement. Find the measure of the angle.
Answer by checkley77(12569)

X+2X+3=180
3X=180-3
3X=177
X=177/3
X=59 DEGREES IS THE SMALLER ANGLE
180-59=121 DEGREES IS THE LARGER ANGLE.
PROOF:
59+2*59+3=180
59+118+3=180
180=180